OpenStudy (anonymous):

PLEASE HELP!!!!!!! Choose the equation of the line passing through the point (3, -2) and parallel to y = 3x - 6.

OpenStudy (jhannybean):

When lines are"parallel" they have the same slope, so if we took the slope from the line given, \(\large y=\color{red}{3}x-6\), we can use that slope, 3, and out points (3,-2) to find a new line using point slope form, \[\large y-y_{1}=m(x-x_{1})\]use the points given and (x1,y1) and the slope of the original line, m=3 \[\large y-(-2)=3(x-3)\]\[\large y+2 =3(x-3)\]\[\large y=3(x-3)-2\] and simplify it by multiplying in the 3 to (x-3) and reducing it. I'll let you finish up.

OpenStudy (anonymous):

eq of line parallel to y=3x-2 is y=3x+c it passes through (3,-2) -2=3*3+c c=-2-9=-11 reqd. eq is y=3x-11

OpenStudy (jhannybean):

I didn't want to give her the solution @surjithayer <_<

OpenStudy (anonymous):

thank you, both of you i kinda understand it now

OpenStudy (jhannybean):

awesome :)